Make New Entry,
Make Followup Entry
User name V. Sulkosky
Log entry time 22:35:40 on May10,2011
Entry number 358877
keyword=High rate during run 4369
It seems one of the cables going to the TS has a bad connection. We
saw a high rate of a 2.6 kHz on the L1A. I will try to isolate the trigger.